# Boosted second moment method in random regular graphs

### v1

@@ 4. Augmented independent sets
In this section we discuss the local improvements that one can make to an independent set with the spatial Markov property. Let us start from a Markovian independent set of density $\alpha$ (see Definition 3.1). If we have two neighboring vertices, then the conditional probability that one has label 0 conditioned on the other having label 0 is equal to
@@
In the introduction we defined a full-zero vertex as a vertex whose label, as well as all its neighbors’ labels, are 0. Clearly, the probability that a vertex is full-zero is
(11) $(1-\alpha)p^{d}.$
Furthermore, the probability that a vertex is an isolated full-zero vertex (that is, a full-zero vertex with no full-zero neighbor) is
(12) $(1-\alpha)p^{d}\big(1-p^{d-1}\big)^{d}.$
Now consider the induced subgraph over all full-zero vertices. We will need that the connected components of this subgraph are finite (almost surely). Note that the conditional probability that a given neighbor of a full-zero vertex is also full-zero is equal to $p^{d-1}$. Therefore, if the condition
$p^{d-1}\leq\frac{1}{d-1}$
holds, then the corresponding Galton–Watson process dies out almost surely, and hence we have finite components. This condition is always comfortably satisfied whenever $\alpha$ satisfies the second moment condition. Therefore, we may assume that the components are finite.
@@
Let $d\geq 3$ be an integer and $\alpha\in(0,1/2)$, and set $p=(1-2\alpha)/(1-\alpha)$. Suppose that the function $f_{d,\alpha}\colon[0,\alpha]\to\mathbb{R}$ attains its maximum at $\alpha^{2}$, and $(d-1)p^{d-1}\leq 1$. Then there exists a typical independent set on $T_{d}$ with density $\hat{\alpha}$ as in (13).
See Table 2 in the introduction for the density $\hat{\alpha}$ of the augmented independent set for small degrees.
One could get a slightly higher density by also computing the probability of components with three vertices, or by considering further local improvements but they would only lead to marginal gains.
@@ 5. Star decompositions
To further demonstrate the strength of our approach, in this section we consider the following problem. Given an integer $d/2<k<d$, can the edges of $\mathbb{G}_{N,d}$ be partitioned into edge-disjoint stars, each containing $k$ edges? We restrict ourselves to those $N$ for which the number of edges ($Nd/2$) is divisible by $k$. If such a partition exists with probability $1-o_{N}(1)$, then we say that $\mathbb{G}_{N,d}$ asymptotically almost surely (a.a.s.) has a $k$-star decomposition.
The case $d=4,k=3$ was settled (affirmatively) in [7]. Later, that method was extended in [6], and a sufficient condition was given for general pairs $d,k$. Both papers used a second moment method and checking the required conditions lead to very technical computations.
@@
Given a positive integer $\hat{d}<d$, we say that a set $A\subset V(G)$ is $\hat{d}-$thin in $G$ if every $v\in V(G)\setminus A$ has at most $\hat{d}$ neighbors in $A$.
@@
If we choose the parameters properly, then conditions (iii) and (iv) actually hold true a.a.s. for random regular graphs. In fact, one can simply check the numerical condition of [13, Cor 3.4] to check (iii) and (iv) in specific cases. We omit the details here as this is not the main focus of the current paper. Instead, we focus on the first two conditions.
@@
Assume that the Markovian independent set of density $\alpha$ is typical, and set $p\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(1-2\alpha)/(1-\alpha)$. For $\hat{d}<d$ there exists a $\hat{d}-$thin invariant independent set $A_{\mathrm{thin}}\subseteq A$ of density at least

### v2

@@ 5. Augmented independent sets
In this section we discuss the local improvements one can make to an independent set with the spatial Markov property. Let us start from a Markovian independent set of density $\alpha$ (see Definition 4.1). If we have two neighboring vertices, then the conditional probability that one has label 0 conditioned on the other having label 0 is equal to
@@
We call a vertex full-zero if its label is 0 and the labels of all its neighbors are also 0. Then, for any vertex $v$ of $T_{d}$, we clearly have
(24) $\mathbb{P}(v\text{ is full-zero})=(1-\alpha)p^{d}.$
We will also need the following conditional probability. If $u$ and $v$ are adjacent vertices, then
(25) $\mathbb{P}(u\text{ is full-zero}\,|\,v\text{ is full-zero})=p^{d-1}.$
It follows that the probability of a vertex being an isolated full-zero vertex (that is, a full-zero vertex with no full-zero neighbor) is
[…]
Now consider the induced subgraph over all full-zero vertices. We will need that the connected components of this subgraph are finite (almost surely), that is, the corresponding Galton–Watson process dies out almost surely. The latter happens if and only if the conditional probability (25) is at most 1/(d1)1/(d-1). Next we check that this condition holds (unless $\alpha$ is too small) so that we may safely assume that the components are finite.
Claim 5.1.
$\text{If }\alpha\geq\frac{\log(d-1)}{d-1}\text{, then }p^{d-1}<\frac{1}{d-1}.$
Proof.
Note that
$p=\frac{1-2\alpha}{1-\alpha}<1-\alpha\leq 1-\frac{\log(d-1)}{d-1}\leq\exp\left(-\frac{\log(d-1)}{d-1}\right).$
Therefore,
$p^{d-1}<\exp\left(-\log(d-1)\right)=\frac{1}{d-1}.$
@@
Let $d\geq 3$ be an integer and $\alpha\in(0,1/2)$, and set $p=(1-2\alpha)/(1-\alpha)$. Consider the function $f_{d,\alpha}\colon[0,\alpha]\to\mathbb{R}$ defined in (21) and suppose that it attains its maximum at $\alpha^{2}$. If $(d-1)p^{d-1}\leq 1$, then there exists a typical independent set on $T_{d}$ with density $\hat{\alpha}$ as in (27).
Setting $\alpha=\alpha^{\mathrm{SM}}_{d}$, we get the boosted bound $\hat{\alpha}=\alpha^{\mathrm{aug}}_{d}$ stated in (4). It can be shown that the condition $(d-1)p^{d-1}\leq 1$ is always satisfied [^14] Why can we assume that the condition $(\ast)\;(d-1)p^{d-1}\leq 1$ (for $\alpha=\alpha^{\mathrm{SM}}_{d}$) holds true? In this paper we only use Corollary 5.3 for producing numerical bounds, in which case it suffices to check $(\ast)$ for the given $d$ (which we did). To show that $(\ast)$ holds true for every $d$, one may combine Theorem 1.3 and Theorem 11.1 to prove that $\alpha^{\mathrm{SM}}_{d}\geq\frac{\log(d-1)}{d-1}$ for sufficiently large $d$. This, in turn, implies $(\ast)$ by Claim 5.1. for this $\alpha$, and hence we get the following corollary.
Corollary 5.3.
For $d\geq 3$ let $\alpha^{\mathrm{aug}}_{d}$ be the density defined in (4). Then $T_{d}$ has a typical independent set of density $\alpha^{\mathrm{aug}}_{d}$. In particular, $\alpha^{\star}_{d}\geq\alpha^{\mathrm{aug}}_{d}$.
Table 1 in the introduction compares the original density $\alpha^{\mathrm{SM}}_{d}$ and the increased density $\alpha^{\mathrm{aug}}_{d}$ for specific degrees.
One could achieve a slightly higher density by also computing the probability of components with three vertices (two of which could be added to the independent set), or by considering further local improvements, but these would only lead to marginal gains.
5.1. The asymptotic gain of the boost
The above local improvement increases the density substantially for small degrees. Asymptotically, however, the gain is only of order $1/d^{2-o(1)}$. More precisely, one can show that
(28) $\alpha^{\mathrm{aug}}_{d}<\alpha^{\mathrm{SM}}_{d}+\frac{(\log d)^{2}}{d^{2}}=\alpha^{\mathrm{SM}}_{d}+\frac{1}{d^{2-o(1)}}.$
From (2) and (6) we have
$\alpha^{\mathrm{SM}}_{d}=\alpha^{\mathrm{FM}}_{d}-\frac{1}{d^{3/2-o(1)}}\quad\text{and}\quad\alpha^{\star}_{d}=\alpha^{\mathrm{FM}}_{d}-\frac{1}{d^{2-o(1)}}.$
It means that (for large $d$) we do not get significantly closer to the true value with the augmented bound $\alpha^{\mathrm{aug}}_{d}$.
We finish this section by sketching an argument for proving (28). As we pointed out in the proof of Claim 5.1, we have $p<1-\alpha$. Moreover, the probability (24) that a vertex is full-zero is clearly an upper bound for the density gain:
$\hat{\alpha}-\alpha\leq(1-\alpha)p^{d}\leq p^{d}<(1-\alpha)^{d}\leq\exp(-d\alpha).$
From Theorem 1.3 and Theorem 11.1(c), one can easily deduce that
$\alpha^{\mathrm{SM}}_{d}\geq\frac{2\log d}{d}-\frac{2\log\log d}{d}.$
Therefore, for $\alpha=\alpha^{\mathrm{SM}}_{d}$ the gain is less than
$\exp\big(-2\log d+2\log\log d\big)=\frac{(\log d)^{2}}{d^{2}}.$
@@ 6. Star decompositions
To further demonstrate the strength of our approach, we consider the following problem. Given an integer $d/2<k<d$, can the edges of $\mathbb{G}_{N,d}$ be partitioned into edge-disjoint stars, each containing $k$ edges? We restrict ourselves to those $N$ for which $Nd/2$ (i.e., the number of edges) is divisible by $k$. If such a partition exists with probability $1-o_{N}(1)$, then we say that $\mathbb{G}_{N,d}$ asymptotically almost surely (a.a.s.) has a $k$-star decomposition.
The case $d=4,k=3$ was settled (affirmatively) in [8]. Later, that method was extended in [7], and a sufficient condition was given for general pairs $d,k$. Both papers used a second moment method and checking the required conditions led to very technical computations.
@@
Given a positive integer $\hat{d}<d$, we say that a set $A\subset V(G)$ is $\hat{d}-$thin in $G$ if every outside vertex $v\in V(G)\setminus A$ has at most $\hat{d}$ neighbors in $A$.
@@
If we choose the parameters properly, then conditions (iii) and (iv) actually hold true a.a.s. for random regular graphs. In fact, one can simply check the numerical condition of [15, Cor 3.4] to check (iii) and (iv) in specific cases. We omit the details here as this is not the main focus of the current paper. Instead, we focus on the first two conditions.
@@
Assume that the Markovian independent set of density $\alpha$ is typical, and let $p\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(1-2\alpha)/(1-\alpha)$. For $\hat{d}<d$ there exists a $\hat{d}-$thin invariant independent set $A_{\mathrm{thin}}\subseteq A$ of density at least